### v1

# Data-Driven Persuasion

Abstract
This paper develops a data-driven approach to Bayesian persuasion. The receiver is privately informed about the prior distribution of the state of the world, the sender knows the receiver’s preferences but does not know the distribution of the state variable, and the sender’s payoffs depend on the receiver’s action but not on the state. Prior to interacting with the receiver, the sender observes the distribution of actions taken by a population of decision makers who share the receiver’s preferences in best response to an unobserved distribution of messages generated by an unknown and potentially heterogeneous signal. The sender views any prior that rationalizes this data as plausible and seeks a signal that maximizes her worst-case payoff against the set of all such distributions. We show positively that the two-state many-action problem has a saddle point and negatively that the two-action many-state problem does not. In the former case, we identify adversarial priors and optimal signals. In the latter, we characterize the set of robustly optimal Blackwell experiments.
JEL classification: D81, D82, D86
Keywords: persuasion, robustness, data-driven mechanism design
1 Introduction
A prosecutor wishes to secure a conviction, the defendant is either innocent or guilty, and the judge is indifferent between type one and type two errors. As in the original formulation of both the example and the general persuasion problem (Kamenica and Gentzkow (2011)), the judge is experienced and knows the prior probability of guilt. Unlike in the standard formulation, the prosecutor is inexperienced and does not. Instead, she knows only that the court convicts $40\%$ of defendants.
What can the prosecutor infer about the unconditional probability of guilt? At one extreme, each of the court’s convictions might have been won by a zealous district attorney who acted to maximize the unconditional probability of conviction. In that case, because half of the convicted defendants and none of the acquitted defendants were guilty, the prior probability of guilt is $20\%$. At the other extreme, each conviction might be due to a reluctant district attorney who sought acquittals for as many defendants as possible. In that case, all of the convicted defendants and half of the acquitted defendants were guilty, and the prior probability of guilt is $70$%.
What should the prosecutor do with this information? In light of her preference for convictions, the robustly optimal communication strategy is to recommend the court convict every guilty defendant and $25\%$ of innocent defendants, as if she were certain that each previous case was brought by a zealous district attorney. This signal is guaranteed to recover a $40\%$ conviction rate and will yield convictions more frequently if a portion of the court’s earlier decisions were in response to fair-minded prosecutions.
This paper proposes a model of robust data-driven persuasion in which the receiver knows the distribution of the state of the world but the sender does not. There are finitely many actions, finitely many states, and the sender’s payoff depends on the former but not on the latter. Prior to interacting with the receiver, the sender observes the distribution of actions taken by a population of decision makers in best response to a distribution of messages generated by an unknown signal. Both the messages generated by these signals and the signals themselves are unobserved and heterogeneous, the decision makers’ preferences are known and homogeneous, and the receiver is a representative member of the population. With full control over what is communicated to the receiver, the sender seeks a signal that maximizes her worst case payoff against the set of prior distributions that rationalize the action data.
Our first task is to identify the set of priors that are consistent with the sender’s observations. In the language of the motivating example, our notion of rationalization is general enough to allow for situations in which the court sometimes hears zealous arguments, sometimes hears reluctant arguments, and sometimes hears arguments that are neither; the court fields recommendations that are more varied than simply “acquit” or “convict”; and an on-the-fence court convicts some defendants and acquits others. While this flexibility might in principle complicate our analysis, we show to the contrary that a given prior rationalizes the action data if and only if it does so via a fixed signal in which each of the district attorneys in the prosecutor’s data recommended convictions and acquittals with homogeneous frequencies and in an incentive compatible manner. With that reduction in hand, we use the supporting hyperplane theorem to characterize the set of rationalizing posteriors and Bayes plausibility (Kamenica and Gentzkow (2011)) to combine rationalizing posteriors into rationalizing priors.
We initiate our analysis of the sender’s problem itself with a formal analysis of the two-action two-state benchmark problem. As in the motivating example, the lowest [^1] In the two-action two-state problem, we call the state in which the receiver prefers the sender’s disfavored action the low state and the state in which he prefers the sender’s favored action the high state. We order priors by the probability they assign to the latter. prior that rationalizes the data and the optimal signal for the counterfactual problem in which the sender knows the state of the world is distributed according to that prior form a saddle point of the sender’s problem. Because improvements to the prior probability of the high state increase the frequency with which the signal recommends the sender’s preferred action and also relax the incentive compatibility constraint associated with that recommendation, their effect is to unambiguously improve the sender’s payoff. The worst-case problem is as if the sender knew the state of the world were distributed according to the lowest rationalizing prior.
Our positive result for the two-state two-action problem problem extends readily to the two-state many-action problem as long as the sender’s payoffs are increasing in the receiver’s action. [^2] In the two-state problem, we label the states low and high arbitrarily and enumerate the receiver’s actions in ascending order according to their payoff in the high state. The substance of our monotonicity assumption lies in requiring that the sender’s utility is increasing in one of the two potential labelings of the receiver’s actions. The lowest prior in the identified set is adversarial; the signal designed for that prior is robustly optimal; and improvements to the prior probability of the high state (i) strictly increase the probability with which the optimal signal sends the higher of its two messages and (ii) weakly increase the receiver’s best response to both of those messages. By monotonicity of the sender’s preferences, increasing the prior improves the sender’s payoff.
In contrast to the two-state many-action problem, our positive result for the two-by-two problem does not extend to the two-action many-state problem under any non-trivial specification of sender and receiver preferences. Unlike in the two-state problem, the set of rationalizing priors is multidimensional and its lower boundary contains a multiplicity of distributions that are not ordered in any relevant sense. While the optimal signal in the known-prior counterfactual recovers the action distribution observed in the data for each one of the boundary priors, any signal that is optimal in one of those counterfactuals is suboptimal in all of the others. No signal is guaranteed to perform as well as the unobserved signal that generated the data and no prior is adversarial.
The non-existence of an adversarial prior in the two-action many-state problem implies that exact knowledge of the true distribution of the state variable is of guaranteed value to the sender. In fact, less information suffices. Formally, we (i) call a Blackwell experiment reliable if its output is guaranteed to induce a saddle point in the sender’s problem and (ii) completely characterize the set of reliable experiments as those that are ordered in terms of the one-dimensional structure that their outcomes impose on the set of rationalizing priors. We interpret reliability as a robust optimality criterion and provide a formal justification for that interpretation in the body; conversely, ordered experiments are those that satisfy a pair of restrictions to the null space of the stochastic matrix that represents the experiment.
Blackwell experiments play a different role in our environment than they do in their traditional setting, wherein the decision maker receives a single message and uses their knowledge of its conditional distribution to update their prior beliefs about the state of the world. In our model, the sender instead observes the entire distribution of messages and uses her knowledge of their conditional distribution to infer restrictions on the prior distribution of the state variable. Within that framework, the simplest class of experiments that satisfy our order criterion are those that pin down the probabilities associated with each of the states in which the receiver strictly prefers the sender’s disfavored action. More generally, while we do characterize the reliability of simple experiments that (i) partition the state space into cells and (ii) reveal the probabilities of each cell in the partition, reliable experiments need not be simple.
The paper is organized as follows. In order to fix ideas, we lay out the model in Section 2 before reviewing the literature in Section 3. In Section 4 we study the identification of the sender’s problem. Next, in Sections 57 we analyze the two-action two-state benchmark problem, the two-state many-action problem, and the two-action many-state problem, respectively. We conclude in Section 8 and present technical material in the Appendix that follows.
2 Model
Given a finite set $X$, we write $\Delta(X)$ for the set of probability distributions on $X$ and $\delta(x)$ for the degenerate distribution at $x$. We interpret elements of $\Delta(X)$ variously as measures or as real vectors, give the real numbers and their products the Euclidean topology, and write $\varepsilon$ to indicate an arbitrarily small strictly positive real number. Finally, given a matrix $M_{m\times n}$, we write $M_{i}\equiv(M_{1i},...,M_{mi})$ for its columns.
2.1 Action, states, and payoffs
Actions are elements $a$ of finite set $A$, states are elements $\omega$ of finite set $\Omega$, and there are at least two actions and at least two states. The receiver’s payoffs are described by utility function $u:A\times\Omega\to\mathbb{R}$, the sender’s payoffs are described by state-independent and injective utility function $v:A\to\mathbb{R}$, and the state of the world is distributed according to prior distribution $\mu\in\Delta(\Omega)$.
2.2 Signals
A signal is a pair $(\pi,S)$, where $S$ is a finite set of messages and the map $\pi:\Omega\to\Delta(S)$ assigns a distribution of messages $\pi(\cdot|\omega)$ to each state $\omega$. Given a prior $\mu\in\Delta(\Omega)$ and a signal $(\pi,S)$, we write
$\mu^{s}(\omega)$ $\equiv\frac{\pi(s|\omega)\mu(\omega)}{\sum_{\omega^{\prime}}\pi(s|\omega^{\prime})\mu(\omega^{\prime})}$
for the posterior induced by message $s$ given prior $\mu\in\Delta(\Omega)$. Given posterior $\mu^{s}\in\Delta(\Omega)$, we write
$a^{*}(\mu^{s})$ $\equiv\underset{a\in A}{\arg\max}\;\sum_{\omega}u(a|\omega)\mu^{s}(\omega),$ $\widehat{a}(\mu^{s})$ $\equiv\underset{a\in a^{*}(\mu^{s})}{\arg\max}\;v(a)$
for the set of receiver-optimal actions and the sender’s preferred element of that set, respectively, noting that $\widehat{a}$ is single-valued per our hypothesis that the sender’s utility $v$ is injective. Ties are broken in favor of the sender and her expected utility from signal $(\pi,S)$ and prior $\mu$ is thus
$V((\pi,S)|\mu)\equiv\sum_{s}\sum_{\omega}v(\widehat{a}(\mu^{s}))\pi(s|\omega)\mu(\omega).$
If $(\pi,S)$ is a direct signal with $S\subset A$, we sometimes interpret the condition $a\in a^{*}(\mu^{a})$ as an incentive compatibility constraint. More generally, in places where the identity of the signal $(\pi,S)$ should be understood, we write $S^{a}(\mu)\equiv\{s\in S|a=\widehat{a}(\mu^{s})\}$ for the set of messages under which the receiver takes action $a$ given prior $\mu$.
2.3 Rationalizing priors
The prior distribution of the state of the world is known to the receiver but not to the sender. Instead, the sender — who does know the receiver’s utility $u$ — observes the distribution of best responses by a population of receivers to some heterogeneous and unobserved signal. She entertains the possibility that the prior might be any that rationalize the data.
Definition 1.
Prior $\mu$ homogeneously rationalizes action distribution $\alpha\in\Delta(A)$ if there exists a signal $(\pi,S)$ and a choice rule $C:S\to\Delta(A)$ such that (i) $a\in a^{*}(\mu^{s})$ for all $a,s$ with $a\in\text{supp}(C(s))$ and (ii) $\sum_{\omega}\sum_{s}C(a|s)\pi(s|\omega)\mu(\omega)=\alpha(a)$ for all $a\in\text{supp}(\alpha)$.
Prior $\mu$ homogeneously rationalizes action distribution $\alpha$ if there exists a fixed signal $(\pi,S)$ and a tiebreaking rule $C$ such that $\alpha$ is the distribution of best responses to $(\pi,S)$ when the state of the world is distributed according to $\mu$ and ties are broken according to $C$.
Definition 2.
Prior $\mu$ rationalizes action distribution $\alpha\in\Delta(A)$ if there exists a finite index $g=1,...,G$, a distribution $\gamma\in\Delta(\{1,...,G\})$, and a collection of action distributions $\alpha^{1},...,\alpha^{G}\in\Delta(A)$ such that (i) $\mu$ homogeneously rationalizes $\alpha^{g}$ for each $g$ and (ii) $\alpha=\sum_{g}\gamma^{g}\alpha^{g}$.
More generally, $\mu$ rationalizes $\alpha$ if there exists a distribution $\gamma$ of signals $(\pi^{g},S^{g})$ such that $\alpha$ is the average distribution of best responses to $(\pi^{g},S^{g})$ when the state of the world is distributed according to $\mu$. We write $\mathcal{P}(\alpha)$ for the set of all such priors.
2.4 The sender’s problem
The sender’s problem
$\max_{(\pi,S)}\;\min_{\mu\in\mathcal{P}(\alpha)}\;V((\pi,S)|\mu)$
is to maximize her guarantee $\min_{\mu\in\mathcal{P}(\alpha)}\;V((\pi,S)|\mu)$ against the set of priors that rationalize the data $\alpha$. Our approach to analyzing this problem is to identify saddle points $((\pi^{*},S^{*}),\mu^{*})$ defined by the inequalities
$\forall(\pi,S)\;V((\pi^{*},S^{*})|\mu^{*})$ $\geq V((\pi,S)|\mu^{*}),$ $\forall\mu\in\mathcal{P}(\alpha)\;V((\pi^{*},S^{*})|\mu)$ $\geq V(\pi^{*},S^{*})|\mu^{*}).$
If the sender’s problem has saddle point $((\pi^{*},S^{*}),\mu^{*})$ then $(\pi^{*},S^{*})$ is robustly optimal and its guarantee is the same as in the counterfactual environment in which the state is known to be distributed according to the adversarial prior $\mu^{*}$. Conversely, if the sender’s problem does not have a saddle point then the value of the sender’s problem is lower than the value of the counterfactual problem in which the sender knows that the prior is $\mu$, for any $\mu$ that rationalizes the data. Accordingly, knowledge of the distribution of the state of the world is of guaranteed value to the sender if and only if her problem lacks a saddle point.
3 Literature review
This paper contributes to a number of partially overlapping segments of the extensive literature on persuasion and information design.
Robust persuasion
We share our emphasis on uncertainty and worst-case solution concepts in persuasion problems with a handful of recent studies. Dworczak and Pavan (2022) study an environment in which there are common prior beliefs about the distribution of the state of the world and the sender is concerned about the presence of unknown background signals. In contrast, our sender has complete control over what is communicated to the receiver.
Elsewhere, Kosterina (2022) studies a problem in which the sender and the receiver have different beliefs about the distribution of the state of the world and the sender is uncertain about the receiver’s beliefs. We differ in two important ways. First, while the uncertainty set in their model is abstract, the uncertainty set in our model it is derived from choice data. Second, while the sender in both models is uncertain about the receiver’s beliefs, the sender in our model believes the receiver knows the true distribution of the state of the world. Accordingly, while our sender is uncertain about both the distribution of messages sent to the receiver and the receiver’s response to those messages, their sender is only uncertain about the latter.
Less closely related is one study in which a regret-minimizing sender is unaware of the receiver’s preferences (Babichenko, Talgam-Cohen, Xu, and Zabarnyi (2022)) and another in which a regret-minimizing sender gradually learns about the state variable via direct observation of its realizations (Castiglioni, Celli, and Gatti (2023)).
Counterfactuals in games with unknown information structures
A number of recent studies conduct counterfactual analyses in environments with unobserved information structures. First, Bergemann, Brooks, and Morris (2022) consider a setting in which an analyst observes the behavior of a group of agents playing a game but does not observe either the distribution of the state variable or the information held by the decision makers. Aside from our focus on individual rather than group behavior and our use of worst-case rather than Bayesian solution concepts, in our model the sender again has full control over the receiver’s information while in theirs the information structure in the observed environment persists in the counterfactual.
Second, Syrgkanis, Tamer, and Ziani (2017) study counterfactual predictions in auctions with unknown information structures. Aside from differences in setting, we focus our counterfactual analysis on worst-case optimal mechanisms, while the authors take their counterfactual mechanisms as given and bound their performance.
Third, Magnolfi and Roncoroni (2023), develop methods for estimation and counterfactual analysis in games with weak assumptions about information and apply those methods to a study of real-world supermarket entry.
Inference about unknown information in decision problems
Our analysis of the identification of the sender’s model is related to a pair of recent papers that study the identification of information in economic settings. First, Doval, Eilat, Liu, and Zhou (2025) consider a single-agent decision problem in which the analyst knows the distribution of the state of the world, the agent’s utility function, and the unconditional distribution of the agent’s actions but does not the agent’s information. By contrast, our sender is ignorant about the distribution of the state of the world, and this uncertainty is the central premise of our paper. Moreover, identification is a preliminary component of our analysis, rather than its central focus.
Second, Libgober (2025) studies an environment in which the analyst knows the posterior beliefs of a group of agents but does not know either the prior distribution of the state world nor the information structure that induced the update from prior to posterior. Because inference about those two objects is the primary objective of that paper, we are differentiated not only by the number of agents but also most importantly by our emphasis on the design of robustly optimal counterfactual information structures.
4 Identification
Before proceeding to our analysis of the sender’s problem, it will be useful to develop an understanding of the identification problem that precedes it. As we have noted in the introduction, our definition of rationalization allows for the action data to be generated in response to heterogeneous signals, signals that send more messages than actions, and complex tiebreaking rules. Our first step is to develop an equivalence with a simpler definition.
Lemma 1.
The following three statements are equivalent:
  1. (i)
    $\mu$ rationalizes $\alpha$;
  2. (ii)
    $\mu$ homogeneously rationalizes $\alpha$;
  3. (iii)
    there exists a signal $(\pi,S)$ such that $S\subset A$, $a\in a^{*}(\mu^{a})$ for all $a\in S$, and $\sum_{\omega}\pi(a|\omega)\mu(\omega)=\alpha(a)$.
Lemma 1 establishes that both of our own criteria are equivalent to rationalization via straightforward signal (Kamenica and Gentzkow (2011)), and the simplicity of these objects facilitates a parsimonious characterization of the identified set via the geometry of the receiver’s payoff vectors. Toward that end, write $\phi:A\to\mathbb{R}^{|\Omega|}$ for the map $a\mapsto(u(a|\omega))_{\omega\in\Omega}$ and $\Phi\subset\mathbb{R}^{n}$ for the free disposal convex hull of its image $\phi(A)$. For each action $a$, write
$\mathcal{N}(a)$$\equiv\{\eta:\Omega\to\mathbb{R}|\sum_{\omega}\eta(\omega)=1,\eta\cdot\phi(a)\geq\eta\cdot x\text{ for all $x\in\Phi$}\},$
$\mathcal{N^{>}}(a)$$\equiv\{\eta:\Omega\to\mathbb{R}|\sum_{\omega}\eta(\omega)=1,\eta\cdot\phi(a)>\eta\cdot x\text{ for all $x\in\Phi$}\}$
for the exterior unit normal vectors that delineate the supporting hyperplanes and strictly supporting hyperplanes to $\Phi$ at $\phi(a)$, respectively, noting that the latter are a subset of the former.
We make two comments about these sets. First, because the payoff set $\Phi$ is unbounded below, the components of $\eta$ are nonnegative for each normal vector $\eta$ in the unit cone $\mathcal{N}(a)$. From that point of view, $\mathcal{N}$ is appropriately interpreted as a correspondence from the set of actions $A$ into the set of probability distributions $\Delta(\Omega)$. Second, with the first point in hand, it is apparent that $\mathcal{N}(a)=\{\mu\in\Delta(\Omega)|a\in a^{*}(\mu)\}$.
Proposition 1.
Prior distribution $\mu$ rationalizes action distribution $\alpha$ if and only if there exists a selection $n$ from $\mathcal{N}$ such that $\mu(\cdot)=\sum_{a\in A}n(\cdot|a)\alpha(a).$
In addition to whatever guidance it might provide on identifying rationalizing priors, Proposition 1 is informative about both when the data is rationalizable at all and when it is rationalizable by multiple priors. In order to state those results, it will be useful to adapt some standard language to our context. Action $a\in A$ is dominated if there exists an action distribution $\gamma\in\Delta(A)$ such that $\sum_{b\in A}u(b|\omega)\gamma(b)>u(a|\omega)$ for all states $\omega$. More generally, action $a\in A$ is redundant if there exists an action distribution $\gamma\in\Delta(A)$ distinct from $\delta(a)$ such that $\sum_{b\in A}u(b|\omega)\gamma(b)\geq u(a|\omega)$ for all states $\omega$. Mathematically, action $a$ is redundant if and only if $\phi(a)$ is not an extreme point of $\Phi$. Economically, $a$ is redundant if it is weakly dominated or a mixture of other actions.
Lemma 2.
First, $\mathcal{N}(a)$ is empty if and only if action $a$ is dominated. Second, $\mathcal{N^{>}}(a)$ is empty if and only if action $a$ is redundant.
While the first part of Lemma 2 falls immediately out of the supporting hyperplane theorem, the second part makes use of the fact that extreme points in finitely generated convex sets have strictly supporting hyperplanes. As a matter of convenience, we assume for the remainder of the paper that none of the receiver’s actions are redundant.
5 The two-state two-action benchmark problem
The two-state two-action problem is a useful benchmark for both the two-state many-action problem and the two-action many-state problem. Label the set of actions $A\equiv\{0,1\}$ to satisfy $v(0)<v(1)$ and the set of states $\Omega\equiv\{l,h\}$ to satisfy $\Delta(l)\equiv u(1|l)-u(0|l)<0<u(1|h)-u(0|l)\equiv\Delta(h),$ noting that the strict inequalities are consistent with our exclusion of redundant actions from the receiver’s decision problem. Further, write
$c\equiv\frac{|\Delta(l)|}{|\Delta(l)|+\Delta(h)}$
for the quantity that satisfies $(1-c)\Delta(l)+c\Delta(h)=0$ and note that action $1$ is optimal for the receiver under posterior belief $\mu^{s}$ if and only if $\mu^{s}(h)\geq c$. Unsurprisingly, the robustly optimal signal is such that this incentive compatibility constraint is exactly satisfied under the lowest prior in the identified set.
Theorem 1.
Suppose there are two states and two actions. Prior $\mu^{*}$ with $\mu^{*}(l)\equiv 1-c\alpha(1),\mu^{*}(h)\equiv c\alpha(1)$ and direct signal $(\pi^{*},S^{*})$ with
$\pi^{*}(a|l)$ $\equiv\frac{1}{\mu^{*}(l)}\begin{cases}1-\alpha(1)+\mu^{*}(h)&a=0\\ \alpha(1)-\mu^{*}(h)&a=1,\end{cases}$ $\pi^{*}(a|h)$$\equiv\begin{cases}0&a=0\\ 1&a=1\end{cases}$
are a saddle point of the sender’s problem with value $\alpha$.
The proof of Theorem 1 is a formalization of our discussion of the motivating example. We direct readers interested in intuition back to the introduction and readers in want of a detailed argument forward to the Appendix.
6 The two-state problem
Label the set of states $\Omega\equiv\{l,h\}$ and the set of actions $A\equiv\{0,...,k\}$. Where convenient, we identify state distributions $\mu$ with the probability $\mu(h)$ they assign to the high state $h$.
Assumption 1.
The sender’s utility function satisfies $v(0)<...<v(k)$ and the receiver’s utility satisfies $u(0|h)<...<u(k|h)$.
The rationalizing priors in the two state problem have a closed form characterization. In light of our exclusion of redundant actions, the receiver’s expected utility functional satisfies a suitable form of quasiconcavity and the bounds $\mathcal{N}(i)=[\xi(i),\xi(i+1)]$ with
$\xi(0)$ $\equiv 0,$ $\xi(i)$ $\equiv\frac{u(i-1|l)-u(i|l)}{u(i-1|l)-u(i|l)+u(i|h)-u(i-1|h)},$ $\xi(k+1)$ $\equiv 1$ (1)
fall immediately out of payoff comparisons between adjacent actions. In turn, Proposition 1 yields the bound
[…]
Define value function $\widehat{v}:[0,1]\to\mathbb{R}$ and its concave envelope $\widehat{V}:[0,1]\to\mathbb{R}$ by
$\widehat{v}(\mu)$ $\equiv v(\widehat{a}(\mu)),$ $\widehat{V}(\mu)$ $\equiv\max\{y|(\mu,y)\in\text{convex hull}(\text{graph}(\widehat{v}))\},$
noting that $\widehat{v}$ is nondecreasing per Assumption 1 and $\widehat{V}$ is well defined because $\widehat{v}$ is upper semicontinuous. Because $(\underline{\mu},\widehat{V}(\underline{\mu}))$ lies on the boundary of convex $\text{convex hull}(\text{graph}(\widehat{v}))$, there exists a pair [^3] The supporting hyperplane theorem improves Caratheodory’s $n+1$ to $n$ on the boundary. of posteriors $p,q$ and a constant $\lambda$ such that
$(\underline{\mu},\widehat{V}(\underline{\mu})))=(1-\lambda)(p,\widehat{v}(p))+\lambda(q,\widehat{v}(q)).$
Consequently, if signal $(\pi,S)$ induces posterior $p$ with probability $1-\lambda$ and posterior $q$ with probability $\lambda$ in the known-prior counterfactual with distribution $\underline{\mu}$ then $V((\pi,S)|\underline{\mu})\geq V((\pi^{\prime},S^{\prime})|\underline{\mu})$ for all signals $(\pi^{\prime},S^{\prime})$. We direct readers interested in further justification of this assertion to Kamenica and Gentzkow (2011) Section II.B.
Theorem 2.
Let Assumption 1 hold, suppose there are two states and three or more actions, and let $p,q,\lambda$ satisfy $(\underline{\mu},\widehat{V}(\underline{\mu}))=(1-\lambda)(p,\widehat{v}(p))+\lambda(q,\widehat{v}(q))$. Prior $\mu^{*}$ with $\mu^{*}(l)\equiv 1-\underline{\mu},\mu^{*}(h)\equiv\underline{\mu}$ and direct signal $(\pi^{*},S^{*})$ with
$\pi^{*}(a|l)$ $\equiv\frac{1}{\mu^{*}(l)}\begin{cases}\mu^{*}(l)-(1-q)\lambda&a=\widehat{a}(p)\\ (1-q)\lambda&a=\widehat{a}(q),\end{cases}$ $\pi^{*}(a|h)$ $\equiv\frac{1}{\mu^{*}(h)}\begin{cases}\mu^{*}(h)-q\lambda&a=\widehat{a}(p)\\ q\lambda&a=\widehat{a}(q)\end{cases}$
are a saddle point of the sender’s problem with value $\widehat{V}(\underline{\mu})$.
The proof of Theorem 2 is a straightforward generalization of Theorem 1, with a few additional steps. Unlike in the two-by-two problem, increasing the prior probability of the high state might change the receiver’s best responses to any of the messages sent by the robustly optimal signal $(\pi^{*},S^{*})$. However, because the sender’s utility is increasing in the receiver’s action and the receiver’s action is increasing in the posterior probability of the high state, any such changes in behavior benefit the sender.
Corollary 1.
Let Assumption 1 hold and suppose there are two states and three or more actions. The value of the sender’s problem $\widehat{V}(\underline{\mu})$ and the value of the data $\sum_{i}v(i)\alpha(i)$ satisfy $V(\underline{\mu})\geq\sum_{i}v(i)\alpha(i)$, with equality if and only if there exists a hyperplane $H$ such that $(\xi(i),\widehat{V}(\xi(i))=(\xi(i),\widehat{v}(\xi(i))\in H$ for all $i$ in the support of $\alpha$.
Corollary 1 follows jointly from Theorem 2 and the observation
$\widehat{V}(\underline{\mu})$ $=\widehat{V}(\sum_{i}\xi(i)\alpha(i))\geq\sum_{i}\widehat{V}(\xi(i))\alpha(i)\geq\sum_{i}\widehat{v}(\xi(i))\alpha(i)=\sum_{i}v(i)\alpha(i),$
where the equalities are definitional, the first inequality is Jensen’s inequality, and the second inequality is by $\widehat{V}\geq\widehat{v}$. The condition for equality reduces to $\widehat{V}(\xi(i))=\widehat{v}(\xi(i))$ on the support of $\alpha$ if $\alpha$ has binary support. Otherwise, if three or more actions appear in the data then violations of the linearity criterion occur for generic specifications of sender and receiver preferences, even if $\widehat{V}$ and $\widehat{v}$ coincide at the appropriate points. We view the availability of guarantees better than the data in typical specifications of our problem as a positive result.
Example 1.
Suppose there are two states and three actions $0,1,2$ with sender utility $v(i)=i$, receiver utility $\phi(0)=(1,-1)$, $\phi(1)=(0,0)$, $\phi(2)=(-2,1)$, and action distribution $\alpha(i)=1/3$.
In Example 1, $\xi(1)=1/2,\xi(2)=2/3,$ prior $\mu^{*}\equiv\underline{\mu}=7/18$ is adversarial and signal $(\pi^{*},S^{*})$ with $S^{*}\equiv\{0,2\}$, $\pi^{*}(2|l)\equiv 7/22,\pi^{*}(2|h)\equiv 1$ is robustly optimal. The sender’s payoff guarantee $V((\pi^{*},S^{*})|\mu^{*})=\widehat{V}(\mu^{*})=7/6$ exceeds the expected value of the data $\sum_{a}v(a)\alpha(a)=1$.
7 The two-action problem
Label the set of actions $A\equiv\{0,1\}$ to satisfy $1\equiv v(1)>v(0)\equiv 0$ and identify action distributions $\alpha$ with the probability $\alpha(1)$ they assign to the sender’s preferred action. Assume the map $\Delta:\Omega\to\mathbb{R}$ defined by $\Delta(\omega)\equiv u(1|\omega)-u(0|\omega)$ is injective, label the states $\Omega\equiv\{\omega_{1},...,\omega_{n}\}$ to satisfy $\Delta(\omega_{1})<...<\Delta(\Omega_{n})$, and recall our exclusion of redundant actions implies $\Delta(\omega_{1})<0<\Delta(\omega_{n})$. For each state $\omega_{i}$, state distribution $\mu$, and signal $(\pi,S)$ write $\Delta_{i}\equiv\Delta(\omega_{i})$, $\mu_{i}\equiv\mu(\omega_{i})$, and $\pi_{i}\equiv\pi(\cdot|\omega_{i})$. Finally, write $L\equiv\{i|\Delta_{i}<0\}$ for the low states, $H\equiv\{i|\Delta_{i}\geq 0\}$ for the high states, and for each pair of states $i\in L,j\in H$ write
$c_{ij}\equiv\frac{|\Delta_{i}|}{|\Delta_{i}|+\Delta_{j}}$
for the quantity that satisfies the indifference condition $(1-c_{ij})\Delta_{i}+c_{ij}\Delta_{j}=0$. It will be helpful to keep in mind that $0\leq c_{ij}\leq 1$, with $c_{ij}<1$ except where $\Delta_{j}=0$.
Our analysis of the two-action problem is organized as follows. First, in Section 7.1 we (i) show the usual revelation principle for persuasion problems does not hold and (ii) we compensate for that failure by developing a complete characterization of optimal signals for the known-prior counterfactual. Next, in Section 7.2 we show that the sender’s problem does not have a saddle point and furthermore that there are no signals guaranteed to recover an action distribution as good or better for the sender than the data. Finally, with that negative result as motivation, we study robustly optimal experimentation in Section 7.3.
7.1 The revelation principle
Unlike in the two-state problem, the restriction to direct signals is at least sometimes with loss in the two-action problem.
Example 2.
There are three states and two actions, the receiver’s utility satisfies $(\Delta_{1},\Delta_{2},\Delta_{3})=(-1,0,1)$, and the data satisfy $\alpha=1/2$.
Consider Example 2. The identified set $\mathcal{P}(\alpha)$ is the set of priors $\mu$ satisfying the system $22\mu_{1}+\mu_{2}\geq 1/2,\mu_{2}+2\mu_{3}\geq 1/$ 2 . Suppose the sender uses direct signal $(\pi,S)$. If $\text{supp}(\pi_{1})$ and $\text{supp}(\pi_{2})$ intersect, then her payoff under prior $(1/2,1/2,0)$ is 00. Alternatively, if $\text{supp}(\pi_{1})$ and $\text{supp}(\pi_{2})$ are disjoint, then her payoff under the prior $(3/4,0,1/4)$ is at most $1/4$. The signal with $\pi_{1}(1)\equiv 0,\pi_{2}(1)\equiv\pi_{3}(1)\equiv 1$ guarantees the sender payoff $1/4$ and is therefore robustly optimal within the class of direct signals.
Suppose the sender instead uses the indirect signal $(\tilde{\pi},\tilde{S})$ with $\tilde{S}\equiv\{s_{1},s_{2},s_{3}\}$ and $\tilde{\pi}_{1}\equiv(5/6,0,1/6),\tilde{\pi}_{2}\equiv(0,1,0),\tilde{\pi}_{3}\equiv(0,0,1)$. If $\mu_{1}\leq 6\mu_{3}$ then the receiver’s best response to messages $s_{2}$ and $s_{3}$ is $1$, the sender’s payoff is $(1/6)\mu_{1}+\mu_{2}+\mu_{3}$, and the worst prior in this region $(3/4,0,1/4)$ yields payoff $3/8$. Alternatively, if $\mu_{1}>6\mu_{3}$ then $1$ is a best response to only $s_{2}$, the sender’s payoff is $\mu_{2}$, and the limiting worst-case prior $(6/10,3/10,1/10)$ in this region yields payoff $3/10$. Accordingly, the indirect signal $(\tilde{\pi},\tilde{S})$ yields a better guarantee $3/10>1/4$ than the best direct signal.
Definition 3.
Signal $(\pi,S)$ is equivalent to direct signal $(\pi^{\prime},S^{\prime})$ for prior $\mu$ if $\sum_{s\in S^{a}(\mu)}\pi_{i}(s)=\pi_{i}^{\prime}(a)$ for all actions $a\in A$ and for all states $\omega_{i}\in\text{supp}(\mu)$.
In light of Example 2, it will be useful to have a complete characterization of the set of optimal signals for the known-prior counterfactual. Toward that end, for each $i=1,...,n$ and each $z\in[0,1]$ define cutoff signal $(\pi^{i,z},S^{i,z})$ by $S^{i,z}\equiv\{0,1\}$ and
$\pi_{j}^{i,z}(0)$$\equiv\begin{cases}1&j<i\\ 1-z&j=i\\ 0&j>i,\end{cases}$ $\pi_{j}^{i,z}(1)$$\equiv\begin{cases}0&j<i\\ z&j=i\\ 1&j>i.\end{cases}$
Prior $\mu$ is eligible if $\sum_{i}\Delta_{i}\mu_{i}<0$. If $\mu$ is indeed eligible, define parameters $(i(\mu),z(\mu))$ and value $\Pi(\mu)$ by
$i(\mu)$ $\equiv\max\{i|\sum_{j\geq i}\Delta_{j}\mu_{j}<0\},$ $z(\mu)$ $\equiv\frac{\sum_{j>i(\mu)}\Delta_{j}\mu_{j}}{|\Delta_{i(\mu)}|\mu_{i(\mu)}},$ $\Pi(\mu)\equiv$ $\sum_{j>i(\mu)}(1+\frac{\Delta_{j}}{|\Delta_{i(\mu)}|})\mu_{j}.$
Otherwise, if $\mu$ is ineligible, define $(i(\mu),z(\mu),\Pi(\mu))\equiv(1,1,1)$. In both cases, we use the shorthand $(\pi^{\mu},S^{\mu})\equiv(\pi^{i(\mu),z(\mu)},S^{i(\mu),z(\mu)})$.
We record a few facts about these signals here. First, the sender’s payoff in the known prior counterfactual satisfies $V((\pi^{\mu},S^{\mu})|\mu)=\Pi(\mu)$. Second, the threshold state $i(\mu)$ belongs to the set of low states $L$ for all priors $\mu$. Third, if $\mu$ is eligible, then $i(\mu)$ belongs to the support of $\mu$ and $z(\mu)<1$. Fourth, and finally, in all cases signal $(\pi^{\mu},S^{\mu}$) is optimal for prior $\mu$ and that optimality is unique up to equivalence.
Proposition 2.
Signal $(\pi^{*},S^{*})$ satisfies $\forall(\pi,S)\;V((\pi^{*},S^{*})|\mu)\geq V((\pi,S)|\mu)$ if and only if $(\pi^{*},S^{*})$ is equivalent to $(\pi^{\mu},S^{\mu})$.
The characterization of optimality in Proposition 2 supports our results on the non-existence of saddle points later in this section.
7.2 Robustly optimal signals
The basic logic of Theorem 1 and Theorem 2 — in which the “lowest” prior is adversarial and its optimal signal is robustly optimal — fails with two actions and three or more states.
Example 3.
There are three states and two actions, the receiver’s utility satisfies $(\Delta_{1},\Delta_{2},\Delta_{3})=(-2,-1,1)$, and the data satisfy $\alpha=1/2$.
First, because $\mu\equiv(1/2,1/4,1/4)$ and prior $\nu\equiv(0,3/4,1/4)$ satisfy $\Pi(\mu)=\Pi(\nu)=\alpha$, both priors rationalize the data. In turn, Proposition 2 implies cutoff signal $(\pi^{\mu},S^{\mu})$ with $(i(\mu),z(\mu))=(1,0)$ is optimal for the known-prior counterfactual in which states are distributed according to $\mu$. However, even though $\nu$ first order stochastically dominates $\mu$, the sender’s disfavored action 00 is the receiver’s unique best response to both messages sent by $(\pi^{\mu},S^{\mu})$ under that prior. Accordingly, $V((\pi^{\mu},S^{\mu})|\nu)=0$ and the signal $(\pi^{\mu},S^{\mu})$ yields no payoff guarantee for the sender. While there might be other signals that do somewhat better for the sender, there are no signals that are guaranteed to an action distribution at least as good as the data.
Theorem 3.
Suppose there are two actions and three or more states. The sender’s problem has a saddle point if and only if $\alpha\in\{0,1\}$.
The proof of Theorem 3 extends the failure of the optimal signal $(\pi^{\mu},S^{\mu})$ in our discussion of the example to the broader class of signals that are equivalent to $(\pi,S)$. Our approach is to construct a pair of priors $\mu,\nu$ with the property that (i) the value of the known-prior counterfactual problem with prior $\mu$ and the value of the known-prior counterfactual problem with prior $\nu$ are both $\alpha$ and (ii) the set of solutions to those two problems do not intersect.
Corollary 2.
Suppose there are two actions and three or more states. There exists a signal $(\pi,S)$ with $\min_{\mu\in\mathcal{P}(\alpha)}V((\pi,S)|\mu)\geq\alpha$ if and only if $\alpha\in\{0,1\}$.
Corollary 2 follows logically from Theorem 3 and the existence of priors with $\Pi(\mu)=\alpha$. We direct readers interested in the construction of such a prior to the theorem’s proof.
7.3 Robustly optimal experiments
The negative result in Theorem 3 suggests that the sender might benefit from learning more about the distribution of the variable, and that is indeed the case. In this section of the paper, we characterize the additional information the sender needs in order to construct a signal that is guaranteed to work at least as well as the signal that generated the data.
Formally, suppose that before choosing a signal $(\pi,S)$ the sender simultaneously observes the action data $\alpha$ and the distribution $\beta$ of output $\sigma$ generated by Blackwell experiment $A_{m\times n}$, where $A_{ti}\equiv\text{Prob}(\sigma=\sigma_{t}|\omega=\omega_{i})$, $\beta_{t}\equiv\text{Prob}(\sigma=\sigma_{t})$, and $\mathcal{Q}(\beta)\equiv\{x\in\Delta(\Omega)|Ax=\beta\}.$ [^4] There are two technicalities to be addressed. First, we restrict attention to outcome distributions $\beta$ with the property that the matrix equation $Ax=\beta$ has solutions. We view this restriction as a misspecification test. Second, every solution to $Ax=\beta$ necessarily satisfies $\sum_{i}x_{i}=1$. Our approach to the non-negativity constraints $x_{i}\geq 0$ is to begin with fully supported distributions $\mu$ and work with perturbations of the form $\mu+d$ for vectors $d$ that are small in every component. In keeping with our analysis of the sender’s problem, our approach is to identify experiments that are guaranteed to induce saddle points in the experimenter’s problem
$\max_{(\pi,S)}\;\min_{\mu\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)}\;V((\pi,S)|\mu).$
Definition 4.
Experiment $A$ is reliable if for all pairs $(\alpha,\beta)$ there exists a signal $(\pi^{*},S^{*})$ and a prior $\mu^{*}\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$ such that for all signals $(\pi,S)$ and for all priors $\mu\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$
$V((\pi^{*},S^{*})|\mu^{*})$ $\geq V((\pi,S)|\mu^{*}),$ $V((\pi^{*},S^{*})|\mu)$ $\geq V((\pi^{*},S^{*})|\mu^{*}).$
If experiment $A$ is reliable, then the experimenter’s problem has a saddle point for every valid specification of the action data $\alpha$ and the experimental data $\beta$. Conversely, if $A$ is not reliable, then there is a specification of the pair $(\alpha,\beta)$ under which that is not the case.
We organize our analysis of optimal experiments into subsection. First, we characterize the set of reliable experiments in Section 7.3.1. Next, in Section 7.3.2 we specialize to simple experiments that partition the state space. Finally, in Section 7.3.3 we consider an alternative timing convention in which the sender chooses the experiment after observing the action data.
7.3.1 Reliable experiments
The properties of the set $\mathcal{Q}(\cdot)$ of priors consistent with the outcome of the sender’s experiment are characterized by the properties of the null space $\mathcal{D}\equiv\{x|Ax=0\}$ of the stochastic matrix $A$. While the distribution of the state of the world is exactly identified if $\mathcal{D}=\{0\}$, we have more generally that prior $\nu$ is consistent with experimental data $A\mu$ if and only if $(\nu-\mu)\in\mathcal{D}$. We make repeated use of the fact that
$\sum_{i}d_{i}=\sum_{i}\sum_{t}A_{ti}d_{i}=\sum_{t}\sum_{i}A_{ti}d_{i}=0$
for all $d\in\mathcal{D}$ and further normalize nonzero directions in $\mathcal{D}$ to satisfy
$\sum_{i\in H}d_{i}$ $\geq 0,$ $\sum_{i\in H}d_{i}=0$ $\implies\sum_{i\in H}\Delta_{i}d_{i}\geq 0,$ $\sum_{i\in H}d_{i}=\sum_{i\in H}\Delta_{i}d_{i}=0$ $\implies d_{\max\{i|d_{i}\neq 0\}}>0.$
Normalizations in hand, the set of reliable experiments has a simple characterization.
Definition 5.
Experiment $A$ is ordered if $\sum_{j\geq i}d_{j}\geq 0$ and $\sum_{j\geq i}\Delta_{j}d_{j}\geq 0$ for all directions $d\in\mathcal{D}$ and for all states $i=1,...,\max L+1$.
Consider the first condition $\sum_{j\geq i}d_{j}\geq 0$ for all $i=1,...,\max L+1$. If $d$ satisfies that condition then for every cutoff signal $(\pi,S)$ movements in the positive direction $\mu+d$ increase the probability with which $(\pi,S)$ recommends the sender’s preferred action to the receiver. Conversely, if $d$ fails the first criterion in Definition 5 then there exists a cutoff signal $(\pi^{i,z},S^{i,z})$ such that movements in the direction $\mu+d$ reduce the probability that the better message is sent. Accordingly, we call the first criterion the frequency criterion.
Consider instead the second condition $\sum_{j\geq i}\Delta_{j}d_{j}\geq 0$ for all $i=1,...,\max L+1$. If $d$ satisfies that condition then for every prior $\mu$ and every cutoff signal movements in the positive direction $\mu+d$ relax the incentive compatibility constraint $\sum_{i}\Delta_{i}(\mu+d)^{1}(\omega_{i})\geq 0$ for the posterior induced by message $1$. Conversely, if $d$ fails the condition then there exists a cutoff signal under which movements in the direction $\mu+d$ tighten that incentive compatibility constraint. We call the second condition the incentive compatibility criterion.
For the purposes of illustration, recall Example 3. There, movements in the direction $d\equiv\nu-\mu=(-1/2,1/2,0)$ satisfy the frequency criterion and the optimal signal $(\pi^{\mu},S^{\mu})$ for prior $\mu$ recommends the sender’s preferred action with higher probability under prior $\mu+\varepsilon d$ than under prior $\mu$. However, because $(\Delta_{1}d_{1},\Delta_{2}d_{2},\Delta_{3}d_{3})=(1,-1/2,0)$, movements in direction $d$ do not satisfy the frequency criterion and the receiver’s best response to message $1$ is action 00. Accordingly, under the interpretation of the sender’s problem as a special case of the experimenter’s problem in which $A$ has rank $1$ and the null space $\mathcal{D}$ of $A$ is the entire set of directions with $\sum_{i}d_{i}=0$, our order criterion fails.
More generally, if both the frequency criterion and the incentive compatibility criterion are satisfied the identified set is one dimensional in the appropriate sense and we recover the positive results of Theorem 1. On the other hand, if one of the criteria fails, then the problem remains multidimensional and the sender is in a position to benefit from additional information about the distribution of the state of the world.
Theorem 4.
Suppose there are two actions and three or more states. Experiment $A$ is reliable if and only if $A$ is ordered.
We develop the proof of Theorem 4 in several steps, with the positive result first as a standalone proposition.
Proposition 3.
Suppose there are two actions and three or more states. If $A$ is ordered then $A$ is reliable.
The basic idea behind Proposition 3 is as follows. Suppose $A$ is ordered and consider any specification of the data $(\alpha,\beta)$. Let prior $\mu\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$ minimize the value function $\Pi(\cdot)$ on the identified set and let $(\pi^{\mu},S^{\mu})$ be the optimal signal for the counterfactual in which the prior is known to be $\mu$.
Consider first movements in the positive direction $\mu+d$. Because $A$ is ordered, these movements (i) increase the probability the high signal is sent per the frequency criterion and (ii) relax the incentive compatibility constraint for that signal per the incentive compatibility criterion. Accordingly, such movements improve the performance of the signal $(\pi^{\mu},S^{\mu})$. Consider instead movements in the negative direction $\mu-d$. As we show, these movements unambiguously degrade the value $\Pi(\cdot)$ of the prior. Because we began by assuming that $\mu$ minimizes that value on the identified set, $\mu-d$ is thus excluded from $\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$. Accordingly, $\mu$ is adversarial and $(\pi^{\mu},S^{\mu})$ is robustly optimal.
Proposition 4.
Suppose there are two actions and three or more states. If $A$ is not ordered then there exists a data set $(\alpha,\beta)$ such that $\min_{\mu\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)}V((\pi,S)|\mu)<\alpha$ for all signals $(\pi,S)$.
The converse of Proposition 3 follows logically from Proposition 4 because the existence of a signal guaranteeing payoff at least $\alpha$ is a necessary condition for the experimenter’s problem to have a saddle point. We develop the result in parts. First, we show that if our order criterion fails then there necessarily exists a direction $d$ under which there is tension between the frequency criterion and the incentive compatibility criterion.
Lemma 3.
Suppose there are two actions and three or more states. A is not ordered if and only if there exists a state $11<i\leq\max L+$ 1 and a direction $d\in\mathcal{D}$ such that either $\sum_{j\geq i}d_{j}>0>\sum_{j\geq i}\Delta_{j}d_{j}$ or $\sum_{j\geq i}\Delta_{j}d_{j}>0>\sum_{j\geq i}d_{j}$.
Second, we show that there exists a prior $\mu$ and corresponding signal $(\pi^{\mu},S^{\mu})$ for which the tension between the frequency criterion and incentive compatibility criterion is relevant.
Lemma 4.
Suppose there are two actions and three or more states. If $i\in L$ and $z\in(0,1)$ then there exists a prior $\mu$ with full support and $(i(\mu),z(\mu))=(i,z)$.
Third, and finally, we confirm that movements in either the positive direction $+d$ or the negative direction $-d$ are consistent with the action data.
Lemma 5.
Suppose there are two actions and three or more states. If $\mu$ has full support and $z(\mu)\in(0,1)$ then for each $d\in\mathcal{D}$ at least one of $\mu+\varepsilon d,\mu-\varepsilon d$ belongs to $\mathcal{P}(\Pi(\mu))$.
Together, Lemmas 34 provide the substance of our proof of Proposition 4. Theorem 4 then follows immediately from the proposition and its predecessor.
Corollary 3.
Suppose there are two actions and three or more states. If experiment $A$ is ordered then for all data sets $(\alpha,\beta)$ there exists a signal $(\pi,S)$ such that $\min_{\mu\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)}V((\pi,S)|\mu)\geq\alpha.$ Conversely, if $A$ is not ordered then there exists a data set $(\alpha,\beta)$ such that $\min_{\mu\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)}\;V((\pi,S)|\mu)<\alpha$ for all signals $(\pi,S)$.
Corollary 3 justifies our reliability criterion on payoff grounds by clarifying that the set of experiments under which the value of the experimenter’s problem with data $(\alpha,\beta)$ is at least $\alpha$ is precisely the set of ordered (and hence reliable) experiments. While the weak inequality in the first part of the suggests that we might refine our criterion by ranking reliable experiments according to the guaranteed value of the information they reveal, no such refinement is possible because the prior $\mu^{\alpha}$ with $\mu_{1}^{\alpha}\equiv 1-c_{1n}\alpha,\mu_{n}^{\alpha}\equiv c_{1n}\alpha$ has value $\Pi(\mu^{\alpha})=\alpha$. Accordingly, the value of the experimenter’s problem with data $(\alpha,A\mu)$ is at most $\alpha$ for every reliable experiment $A$.
7.3.2 Simple experiments
The interpretation of our characterization of reliability is complicated by the need to relate the physical description of the sender’s experiment $A$ to the properties of its null space $\mathcal{D}$. Aside from whatever analytical burden this intermediate step might impose, it renders our order criteria at least somewhat opaque. Toward a remedy, one example of a natural and easy-to-interpret class of experiments are those that pin down the probabilities of some collection of subsets of the state space $\Omega$. In this section, we use Theorem 4 to determine whether or not they are reliable.
Definition 6.
Experiment $A$ is simple if there exists a partition $\{\Sigma_{1},...,\Sigma_{k}\}$ of $\Omega$ with the property that for each consistent outcome $\beta$ there exists a probability distribution $\mu^{\beta}\in\Delta(\{1,...,k\})$ such that $\mu\in\mathcal{Q}(\beta)\iff(\mu(\Sigma_{1}),...,\mu(\Sigma_{k}))=(\mu_{1}^{\beta},...,\mu_{k}^{\beta}).$
If the columns of $A$ are the canonical basis for $\mathbb{R}^{k}$ then $A$ is simple. More broadly, $A$ is simple if and only if its columns modulo identity are linearly independent.
Proposition 5.
Experiment $A_{m\times n}$ is simple if and only if there exists a partition $\{\Sigma_{1},...,\Sigma_{k}\}$ of $\Omega$ and a matrix $B_{m\times k}$ such that (i) $A_{i}=B_{t}$ for all $i\in\Sigma_{t}$ and (ii) $B$ has rank $k$.
In order to state our main result for this section, it will be useful to introduce some terminology. If simple experiment $A$ induces partition $\{\Sigma_{1},...,\Sigma_{k}\}$ then we say (i) $A$ separates state $i$ if there exists an index $t$ with $\Sigma_{t}=\{i\}$ and (ii) $A$ pools states $E\subset\Omega$ if there exists an index $t$ with $\Sigma_{t}=E$.
Theorem 5.
Suppose there are two actions and three or more states. Simple experiment $A$ is reliable if and only if (i) $A$ separates every low state $i$ or (ii) $A$ pools one low state $i$ with one high state $j$ and separates all other states.
Theorem 5 follows jointly from Theorem 4 and a straightforward analysis of the structure of the null space $\mathcal{D}$ associated with simple experiments. By way of interpretation, experiments $A$ that separate each of the negative states are reliable because the prior $\mu$ that minimizes the sum $\sum_{i\in H}\Delta_{i}\mu_{i}$ over the identified set $\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$ is apparently adversarial. Otherwise, if $A$ pools one low state $i$ with one high state $j$ and separates all other states, then the prior that maximizes the probability of $i$ over $\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$ is itself adversarial.
7.3.3 Sequential experiments
In formulating the experimenter’s problem, we have implicitly assumed that the sender must decide on an experiment $A$ before she observes the action data $\alpha$. While this timing assumption strengthens the sufficiency result in Proposition 3, it also weakens the necessity result in Proposition 4. In this section, we show that our results persist if the sender instead observes the action data before she settles on an experiment.
Definition 7.
Given action distribution $\alpha$, experiment $A$ is sequentially reliable if for all outcomes $\beta$ there exists a signal $(\pi^{*},S^{*})$ and a prior $\mu^{*}\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$ such that for all signals $(\pi,S)$ and all priors $\mu\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$
$V((\pi^{*},S^{*})|\mu^{*})$ $\geq V((\pi,S)|\mu^{*}),$ $V((\pi^{*},S^{*})|\mu)$ $\geq V((\pi^{*},S^{*})|\mu^{*}).$
Our order criteria characterize sequential reliability under two additional assumptions. First, we assume for expositional convenience that the action distribution is interior.
Assumption 2.
The action distribution $\alpha$ satisfies $\alpha\in(0,1).$
Second, and substantively, we assume that the state space satisfies a straightforward richness condition.
Assumption 3.
There exists a state $i$ with $\Delta_{i}=0$.
The existence of a state in which the receiver is indifferent between his two actions broadens the set of rationalizations of the data considerably. We imagine a decision problem in which the state variable $\omega$ is ordered with many potential values and the indifference state $\omega_{i}$ is a minimum acceptable threshold for a positive outcome. Examples of such thresholds abound in binary decision problems, wherein we see hard cutoffs in “pass/fail” assessments, 95% thresholds for “statistical significance”, minimum test scores, grades, or class ranks in admissions decisions and hiring decisions, and so on.
Lemma 6.
Suppose there are two actions and three or more states and let Assumptions 23 hold. If $i\in L$ and $z\in(0,1)$ then there exists a prior $\mu$ with full support, $(i(\mu),z(\mu))=(i,z)$, and $\Pi(\mu)=\alpha$.
Lemma 4 is an analogue to our earlier Lemma 4, here for fixed action distributions. Our characterization of optimal experiments then obtains under the same line of argument as before.
Theorem 6.
Suppose there are two actions and three or more states and let Assumptions 23 hold. Experiment $A$ is sequentially reliable if and only if $A$ is ordered.
In the Appendix, we briefly outline how Lemma 4 facilitates an adaptation of the proof of Theorem 4 to the sequential timing convention considered in this section.
Corollary 4.
Suppose there are two actions and three or more states and let Assumptions 23 hold. If experiment $A$ is ordered then for all experimental outcomes $\beta$ there exists a signal $(\pi,S)$ such that $\min_{\mu\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)}\;V((\pi,S)|\mu)\geq\alpha.$Conversely, if experiment $A$ is not ordered then there exists an experimental outcome $\beta$ such that $\min_{\mu\in\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)}\;V((\pi,S)|\mu)<\alpha$ for all signals $(\pi,S)$.
In Corollary 4, we provide an interpretation of sequential reliability as a dynamic optimality criterion for the version of our model in which the action data are revealed before the sender chooses an experiment. The justification for its statements are the same as the justifications for their analogues in Corollary 3.
8 Conclusions
This paper proposes a data-driven model of robust information provision to a single receiver. We first characterize the set of priors that are consistent with observed behavior and second identify robustly optimal signals for a variety of specifications of the sender’s problem. In light of our restriction to models with either two actions or two states, we leave open for future work the analysis of the many-state many-action problem. While our expectation is that the results for those models are more closely related to our analysis of the two-action problem than the two-state problem, it is unclear whether or not our characterization of optimal experiments might generalize in some useful way to that setting.
?refname?
  • Babichenko et al. (2022) Babichenko, Y., I. Talgam-Cohen, H. Xu, and K. Zabarnyi (2022). Regret-minimizing bayesian persuasion. Games and Economic Behavior 136, 226–248.
  • Bergemann et al. (2022) Bergemann, D., B. Brooks, and S. Morris (2022). Counterfactuals with latent information. American Economic Review 112(1), 343–368.
  • Bertsimas and Tsitsiklis (1997) Bertsimas, D. and J. N. Tsitsiklis (1997). Introduction to Linear Optimization. Athena Scientific. Theorem 2.3.
  • Castiglioni et al. (2023) Castiglioni, M., A. Celli, and N. Gatti (2023). Learning to persuade on the fly: Robustness against ignorance. Operations Research 71(4), 1173–1194.
  • Doval et al. (2025) Doval, L., R. Eilat, T. Liu, and Y. Zhou (2025). Revealed information. Unpublished working paper.
  • Dworczak and Pavan (2022) Dworczak, P. and A. Pavan (2022). Preparing for the worst but hoping for the best: Robust bayesian persuasion. Econometrica 90(5), 2017–2051.
  • Kamenica and Gentzkow (2011) Kamenica, E. and M. Gentzkow (2011). Bayesian persuasion. American Economic Review 101(6), 2590–2615.
  • Kosterina (2022) Kosterina, S. (2022). Persuasion with unknown beliefs. Theoretical Economics 17(4), 1537–1572.
  • Libgober (2025) Libgober, J. (2025). Identifying wisdom (of the crowd): A regression approach. Journal of Political Economy Microeconomics. Forthcoming.
  • Magnolfi and Roncoroni (2023) Magnolfi, L. and C. Roncoroni (2023). Estimation of discrete games with weak assumptions on information. The Review of Economic Studies 90(4), 2006–2041.
  • Rockafellar (1970) Rockafellar, R. T. (1970). Convex Analysis. Princeton University Press. Theorem 19.1.
  • Syrgkanis et al. (2017) Syrgkanis, V., E. Tamer, and J. Ziani (2017). Inference on auctions with weak assumptions on information. Unpublished working paper.
?appendixname? A Appendix
A.1 Section 4
Proof of Lemma 1 .
First, we show (ii) implies (iii). Second, we use the first part to simplify our argument that (i) implies (ii). Because (iii) vacuously implies (i), those two demonstrations are sufficient to establish the claimed equivalence.
Proceeding with the first claim, suppose signal prior $\mu$, signal $(\pi,S)$, and incentive compatible choice rule $C:S\to\Delta(A)$ are such that (i) $\sum_{\omega}\sum_{s}C(a|s)\pi(s|\omega)\mu(\omega)=\alpha(a)$ for all $a$ and (ii) $a\in a^{*}(\mu^{s})$ for all $a\in\text{supp}(C(s))$ for all $s$. Consider the signal$(\pi^{\prime},S^{\prime})$ with $S^{\prime}\equiv\text{supp}(\alpha)$ and $\pi^{\prime}(a|\omega)\equiv\sum_{s}C(a|s)\pi(s|\omega)$. For each $a\in S^{\prime}$ define $S^{a}\equiv\{s\in S|a\in\text{supp}(C(s))\}$ and for each $s\in S^{a}$ define
$\lambda(s|a)\equiv\frac{C(a|s)\sum_{\omega}\pi(s|\omega)\mu(\omega)}{\sum_{\omega}\sum_{s\in S^{a}}C(a|s)\pi(s|\omega)\mu(\omega)}.$
First, for each $a\in S^{\prime}$ we have $\lambda(\cdot|a)\in\Delta(S^{a})$ and $\sum_{s\in S^{a}}\lambda(s|a)\mu^{s}=\mu^{a}$, where we write $\mu^{s}$ for the posteriors induced by $(\pi,S)$ and $\mu^{a}$ for the posteriors induced by $(\pi^{\prime},S^{\prime})$. Second, for each $a\in S^{\prime}$ the set $\{\mu|a\in a^{*}(\mu)\}$ is (i) convex and (ii) contains posterior $\mu^{s}$ for each $s\in S^{a}$ by hypothesis. Together, these two facts imply $\mu^{a}\in\{\mu|a\in a^{*}(\mu)\}$ and thus $a\in a^{*}(\mu^{a})$, as required. Third, and finally, we have as claimed $\sum_{\omega}\pi^{\prime}(a|\omega)\mu(\omega)=\sum_{\omega}\sum_{s\in S}\varphi(a|s)\pi(s|\omega)\mu(\omega)=\alpha(a).$
Proceeding with the second claim, suppose $\mu$ rationalizes $\alpha$. By definition, there exists an index $g=1,...,G$, distribution $\gamma\in\Delta(\{1,...,G\})$, and action distributions $\alpha^{1},...,\alpha^{G}$ such that $\mu$ homogeneously rationalizes $\alpha^{g}$ for each $g$ and $\alpha=\sum_{g}\gamma^{g}\alpha^{g}$. Because we have already shown (ii) implies (iii), for each $g$ there exists a signal $(\pi^{g},S^{g})$ such that $S^{g}\subset A$, $a\in a^{*}((\mu^{g})^{a})$ for all $a\in S^{g}$, and $\sum_{\omega}\pi^{g}(a|\omega)\mu(\omega)=\alpha^{g}(a)$ for all $a$ in the support of $\alpha^{g}$, where we write $(\mu^{g})^{a}$ for the posterior induced by signal $(\pi^{g},S^{g})$ and message $a\in S^{g}$. Define signal $(\pi,S)$ by $S\equiv\text{supp}(\alpha)$ and $\pi(a|\omega)\equiv\sum_{g}\gamma^{g}\pi^{g}(a|\omega)$. For each $a\in S$ consider the convex combination $\lambda(\cdot|a)\in\Delta(\{1,...,G\})$ defined by
$\lambda(g|a)\equiv\frac{\gamma^{g}\alpha^{g}(a)}{\alpha(a)}.$
The family of posteriors
$(\mu^{g})^{a}(\omega)\equiv\frac{\pi^{g}(s|\omega)\mu(\omega)}{\sum_{\omega^{\prime}}\pi^{g}(s|\omega^{\prime})\mu(\omega^{\prime})}$
induced by signal $(\pi^{g},S^{g})$ and message $a$ and the posterior
$\mu^{a}(\omega)\equiv\frac{\pi(a|\omega)\mu(\omega)}{\sum_{\omega^{\prime}}\pi(a|\omega^{\prime})\mu(\omega^{\prime})}$
induced by signal $(\pi,S)$ and message $a$ jointly satisfy $\mu^{a}=\sum_{g}\lambda(g|a)(\mu^{g})^{s}$ for each $a$. Because $a\in a^{*}((\mu^{g})^{a})$ by hypothesis and the set $\{\mu|a\in a^{*}(\mu)\}$ is convex, we conclude $a\in a^{*}(\mu^{a})$. In turn, because $\sum_{\omega}\pi(a|\omega)\mu(\omega)=\alpha(a)$, $\mu$ homogeneously rationalizes $\alpha$ via signal $(\pi,S)$. ∎
Proof of Proposition 1.
Suppose there exists a selection $n$ from $\mathcal{N}$ such that $\mu=\sum_{a}n(\cdot|a)\alpha(a)$. Consider signal $(\pi,S)$ defined by $S\equiv\text{supp}(A)$ and
$\pi(a|\omega)\equiv\frac{\alpha(a)n(\omega|a)}{\mu(\omega)},$
with $\pi(\cdot|\omega)$ assigned arbitrarily for states $\omega$ not in $\text{supp}(\mu)$. Direct calculation yields $\mu^{a}(\omega)=n(\omega|a),\sum_{\omega}\pi(a|\omega)\mu(\omega)=\alpha(a)$ and thus Lemma 1 implies $\mu$ rationalizes $\alpha$.
Conversely, suppose $\mu$ rationalizes $\alpha$. Again per Lemma 1, there exists a signal $(\pi,S)$ such that (i) $S\subset A$; (ii) $a\in a^{*}(\mu^{a})$ for all $a\in S$; and (iii) $\sum_{\omega}\pi(a|\omega)\mu(\omega)=\alpha(a)$. First, because $a\in a^{*}(\mu^{a})$ and $\mathcal{N}(a)=\{\mu|a\in a^{*}(\mu)\}$ we have $\mu^{a}\in\mathcal{N}(a)$. Accordingly, $a\mapsto\mu^{a}$ is a selection from $\mathcal{N}$. Second, Bayes rule implies
$\sum_{a}\mu^{a}(\omega)\alpha(a)=\sum_{a\in A}\Big{(}\frac{\pi(a|\omega)\mu(\omega)}{\sum_{\omega^{\prime}}\pi(a|\omega^{\prime})\mu(\omega^{\prime})}\Big{)}\alpha(a)=\sum_{a\in A}\pi(a|\omega)\mu(\omega)=\mu(\omega)$
for each state $\omega$ and hence that $\mu=\sum_{a}\mu^{a}\alpha(a)$. ∎
Proof of Lemma 2.
First, if action $a$ is dominated then payoff vector $\phi(a)$ belongs to the interior of $\Phi$ and thus $\mathcal{N}(a)$ is empty. Conversely, if $a$ is not dominated then $\phi(a)$ belongs to the boundary of $\Phi$ and the supporting hyperplane theorem implies $\mathcal{N}(a)$ is nonempty.
Second, action $a$ is neither dominated nor redundant if and only if payoff vector $\phi(a)$ is an extreme point of $\Phi$. We claim more strongly that $\phi(a)$ exposed. [^5] Point $\phi(a)$ is an extreme point of $\Phi$ if $\phi(a)$ is not a convex combination of other points in $\Phi$. More specifically, $\phi(a)$ is an exposed point of $\Phi$ if there exists a hyperplane that (i) supports $\Phi$ at $\phi(a)$ and (ii) intersects $\Phi$ only at $\phi(a)$. This fact follows readily from two well known results. First, because $\Phi$ is a finitely generated convex set in the sense of Rockafellar (1970), Theorem 19.1 of that text implies that $\Phi$ is the intersection of finitely many closed half spaces. Second, because $\Phi$ is therefore a polyhedron in the sense of Bertsimas and Tsitsiklis (1997), Theorem 2.3 of that text implies that $\phi(a)$ is indeed exposed. By definition, then, exists a vector $\eta\in\mathcal{N^{>}}(a)$ such that $\eta\cdot\phi(a)>\eta\cdot x$ for all $x\in\Phi$. ∎
A.2 Proof of Theorem 1
First, it is apparent that $V((\pi^{*},S^{*})|\mu^{*})\geq V((\pi,S)|\mu^{*})$ for all signals $(\pi,S)$. [^6] We direct the interested reader to Proposition 2 for a self-contained proof. Second, Proposition 1 implies $\mu(h)\geq\mu^{*}(h)$ for all $\mu\in\mathcal{P}(\alpha)$. Third, and finally, we have
$\mu(h)\geq\mu^{*}(h)\implies$ $\sum_{\omega}\pi^{*}(1|\omega)(\mu(\omega)-\mu^{*}(\omega)),\sum_{\omega}\Delta(\omega)(\mu^{1}(\omega)-(\mu^{*})^{1}(\omega))\geq 0$
$\implies$ $V((\pi^{*},S^{*})|\mu)\geq V((\pi^{*},S^{*})|\mu^{*}).$
Accordingly, $V((\pi^{*},S^{*})|\mu)\geq V((\pi^{*},S^{*})|\mu^{*})=\alpha(1)$ for all rationalizing priors $\mu$.
A.3 Proof of Theorem 2
Where convenient, we identify priors and posteriors with the probability they assign to the high state $h$. Proceeding, per the discussion in the body we have $V((\pi^{*},S^{*})|\mu^{*})\geq V((\pi,S)|\mu^{*})$ for all signals $(\pi,S)$. Let $\mu$ be any rationalizing prior. We claim $V((\pi^{*},S^{*})|\mu)\geq V((\pi^{*},S^{*})|\mu^{*})$.
First, (2) implies $\mu\geq\mu^{*}$ and thus $\mu^{s}\geq(\mu^{*})^{s}$ for all messages $s\in S^{*}$. In turn, because $\widehat{v}$ is nondecreasing, we have $\widehat{v}(\mu^{s})\geq\widehat{v}((\mu^{*})^{s})$ for all $s\in S^{*}$. Suppose without loss of generality $p\geq q$. Again because $\widehat{v}$ is nondecreasing, we have $\widehat{v}(p)\geq\widehat{v}(q)$. Furthermore, because $p\geq q\implies p\geq\mu^{*}\geq q$, direct calculation yields $\pi^{*}(\widehat{a}(p)|h)\geq\pi^{*}(\widehat{a}(p)|l)$. Accordingly, $\sum_{\omega}\pi^{*}(\widehat{a}(p)|\omega)\mu(\omega)\geq\sum_{\omega}\pi^{*}(\widehat{a}(p)|\omega)\mu^{*}(\omega).$ Altogether we have
$V((\pi^{*},S^{*})|\mu)$ $=\sum_{\omega}\pi^{*}(\widehat{a}(p)|\omega)\mu(\omega)\widehat{v}(\mu^{\hat{a}(p)}))+\sum_{\omega}\pi^{*}(\widehat{a}(q)|\omega)\mu(\omega)\widehat{v}(\mu^{\hat{a}(q)}))$
$\geq\sum_{\omega}\pi^{*}(\widehat{a}(p)|\omega)\mu(\omega)\widehat{v}(p)+\sum_{\omega}\pi^{*}(\widehat{a}(q)|\omega)\mu(\omega)\widehat{v}(q)$
$=\sum_{\omega}\pi^{*}(\widehat{a}(p)|\omega)\mu(\omega)(\widehat{v}(p)-\widehat{v}(q))+\widehat{v}(q)$
$\geq\sum_{\omega}\pi^{*}(\widehat{a}(p)|\omega)\mu^{*}(\omega)(\widehat{v}(p)-\widehat{v}(q))+\widehat{v}(q)=V((\pi^{*},S^{*})|\mu^{*}),$
where the first inequality follows jointly from (i) $\widehat{v}(\mu^{s})\geq\widehat{v}((\mu^{*})^{s})$ for all $s\in S^{*}$, (ii) $(\mu^{*})^{\hat{a}(p)}=p,$ and (iii) $(\mu^{*})^{\hat{a}(q)}=q$; and the second follows jointly from $\widehat{v}(p)\geq\widehat{v}(q)$ and $\sum_{\omega}\pi^{*}(\widehat{a}(p)|\omega)\mu(\omega)\geq\sum_{\omega}\pi^{*}(\widehat{a}(p)|\omega)\mu^{*}(\omega).$ The case with $p<q$ is symmetric.
A.4 Proof of Proposition 2
Let $\mu$ be any prior and $(\pi,S)$ be any signal. First, if $\mu$ is ineligible then apparently $(\pi,S)$ is optimal if and only if $(\pi,S)$ is equivalent to $(i(\mu),z(\mu))=(1,1)$. Second, if $\mu$ is eligible and $\text{supp}(\mu)\cap H$ is empty, then every signal is both vacuously optimal and equivalent to $(\pi^{\mu},S^{\mu})$ with $(i(\mu),z(\mu))=(\max\text{supp}(\mu),0)$. Third, if $\mu$ is eligible and $\text{supp}(\mu)\cap H=\min H$ with $\Delta_{\min H}=0$, then $(\pi,S)$ is optimal if and only if $(\pi,S)$ is equivalent to $(\pi^{\mu},S^{\mu})$ with $(i(\mu),z(\mu))=(\max(\text{supp}(\mu)\cap L),0)$. Suppose to the contrary $\sum_{i}\Delta_{i}\mu_{i}<0<\sum_{i\in H}\Delta_{i}\mu_{i}$ and define
$\nu^{0}$ $\equiv\frac{1}{1-V((\pi,S)|\mu)}\sum_{s\in S^{0}(\mu)}\sum_{i}\pi_{i}(s)\mu_{i}\mu^{s},$ $\nu^{1}$ $\equiv\frac{1}{V((\pi,S)|\mu)}\sum_{s\in S^{1}(\mu)}\sum_{i}\pi_{i}(s)\mu_{i}\mu^{s}.$
There are two cases to consider. If $\max\text{supp}(\nu^{0})>\min\text{supp}(\nu^{1})$ then $(\pi,S)$ is apparently not equivalent to $(\pi^{\mu},S^{\mu})$. We claim further that $(\pi,S)$ is not optimal for $\mu$. Define
$i$ $\equiv\max\{j|\sum_{l\geq j}\mu_{l}\geq V((\pi,S)|\mu)\},$ $z$ $\equiv\frac{V((\pi,S)|\mu)-\sum_{j>i}\mu_{j}}{\mu_{i}},$
consider the cutoff signal $(\pi^{i,z},S^{i,z})$, and write $\tilde{\nu}^{1}$ for the posterior induced by its signal $1$. By construction, $\tilde{\nu}^{1}$ is both distinct from $\nu^{1}$ and first order stochastically dominates $\nu^{1}$. Because $\Delta_{i}$ is strictly increasing in $i$, we therefore have $\sum_{i}\Delta_{i}\tilde{\nu}_{i}^{1}>\sum_{i}\Delta_{i}\nu_{i}^{1}\geq 0$. If $z<1$, then $V((\pi^{i,z+\varepsilon},S^{i,z+\varepsilon})|\mu)>V((\pi,S)|\mu)$. Otherwise, if $z=1$, then $V((\pi^{j,\varepsilon},S^{j,\varepsilon})|\mu)>V((\pi,S)|\mu)$ for $j\equiv\max\{k|k<i,k\in\text{supp}(\mu)\}$. In either case, $(\pi,S)$ is not an optimal signal for prior $\mu$.
Alternatively, if $\min\text{supp}(\nu^{1})\geq\max\text{supp}(\nu^{0})$ then there exists a cutoff signal $(\pi^{i,z},S^{i,z})$ that induces posteriors $\nu^{0},\nu^{1}$ and hence satisfies $V((\pi,S)|\mu)=V((\pi^{i,z},S^{i,z})|\mu)$. First, if $V((\pi^{i,z},S^{i,z})|\mu)>V((\pi^{\mu},S^{\mu})|\mu)$ then $\sum_{i}\Delta_{i}\nu_{i}^{1}<0$, in contradiction to the definition of $S^{1}(\mu)$. Accordingly, $(\pi^{\mu},S^{\mu})$ is optimal for prior $\mu$. Second, if $V((\pi^{i,z},S^{i,z})|\mu)=V((\pi^{\mu},S^{\mu})|\mu)$ then $(\pi^{\mu},S^{\mu})$ necessarily induces posteriors $\nu^{0},\nu^{1}$. Consequently,
$\nu_{j}^{1}$ $=\frac{1}{V((\pi,S)|\mu)}\sum_{s\in S^{1}(\mu)}\pi_{j}(s)\mu_{j}=\frac{1}{V((\pi^{\mu},S^{\mu})|\mu)}\pi_{j}^{\mu}(1)\mu_{j}$
for all $j$. Because $V((\pi,S)|\mu)=V((\pi^{i,z},S^{i,z})|\mu)=V((\pi^{\mu},S^{\mu})|\mu)$, we have for all states $j$ $\sum_{s\in S^{1}(\mu)}\pi_{j}(s)\mu_{j}=\pi_{j}^{\mu}(1)\mu_{j}$. We conclude that if $(\pi,S)$ is optimal for prior $\mu$ then $(\pi,S)$ is equivalent to $(\pi^{\mu},S^{\mu})$.
A.5 Proof of Theorem 3
With regards to the two trivial cases, if $\alpha=0$ then $\mu\equiv\delta(1)\in\mathcal{P}(\alpha)$. Apparently, $(i(\mu),z(\mu),\Pi(\mu))=(1,0,0)$ and $((\pi^{\mu},S^{\mu}),\mu)$ are a saddle point of the sender’s problem. Otherwise, if $\alpha=1$ then no prior $\mu$ in the identified set $\mathcal{P}(\alpha)$ is eligible. In turn, we have $(i(\mu),z(\mu),\Pi(\mu))=(1,1,1)$ by definition and $((\pi^{\mu},S^{\mu}),\mu)$ are a saddle point for every $\mu\in\mathcal{P}(\alpha)$. Alternatively, suppose $\alpha\in(0,1)$. Our approach is to show that there does not exist a signal with guarantee at least $\alpha$. There are two cases to consider. First, suppose $|L|=1$ and consider prior $\mu^{\lambda}$ with $\mu_{1}^{\lambda}\equiv 1-\lambda c_{12}\alpha-(1-\lambda)c_{13}\alpha,\mu_{2}^{\lambda}\equiv\lambda c_{12}\alpha,\mu_{3}^{\lambda}\equiv(1-\lambda)c_{13}\alpha$. For all $\lambda$, direct calculation yields
$i(\mu^{\lambda})$ $=1,$ $z(\mu^{\lambda})$ $=\frac{\alpha-\lambda c_{12}\alpha-(1-\lambda)c_{13}\alpha}{1-\lambda c_{12}\alpha-(1-\lambda)c_{13}\alpha},$ $\Pi(\mu^{\lambda})$ $=\alpha.$
In turn, Proposition 2 implies for all $\lambda\in(0,1)$
$V((\pi,S)|\mu^{\lambda})=\alpha\implies\sum_{s\in S^{1}(\mu^{\lambda})}\pi_{1}(s)=z(\mu^{\lambda}),\sum_{s\in S^{1}(\mu^{\lambda})}\pi_{2}(s)=\sum_{s\in S^{1}(\mu^{\lambda})}\pi_{3}(s)=1.$
The necessary conditions for $\pi_{2}$ and $\pi_{3}$ imply the set of incentive compatible messages $S^{1}(\mu^{\lambda})=\text{supp}(\pi_{2})\cup\text{supp}(\pi_{3})$ for prior $\mu^{\lambda}$ does not vary with $\lambda$. On the other hand, because the map $\lambda\mapsto z(\mu^{\lambda})$ is injective, the necessary condition for $\pi_{1}$ implies the set $S^{1}(\mu^{\lambda})$ does vary with $\lambda$. Accordingly, any particular signal $(\pi,S)$ satisfies all three necessary conditions for at most one $\lambda$ in $(0,1)$.
Alternatively, suppose $|L|>1$ and consider prior $\mu$ with $\mu_{1}\equiv 1-c_{1n}\alpha,\mu_{n}\equiv c_{1n}\alpha$ and prior $\nu$ with $\nu_{1}\equiv 1-\alpha,\nu_{2}\equiv\alpha-c_{2n}\alpha,\nu_{n}\equiv c_{2n}\alpha$. It can quickly be verified
$(i(\mu),z(\mu),\Pi(\mu))$ $=(1,\frac{\alpha-c_{1n}\alpha}{1-c_{1n}\alpha},\alpha),$ $(i(\nu),z(\nu),\Pi(\nu))$ $=(1,0,\alpha).$
In turn, Proposition 2 yields the apparently inconsistent conditions
$V((\pi,S)|\mu)\geq\alpha\implies$ $\sum_{s\in S^{1}(\mu)}\pi_{1}(s)>0,\sum_{s\in S^{1}(\mu)}\pi_{n}(s)=1\implies\text{supp}(\pi_{1})\cap\text{supp}(\pi_{n})\neq\emptyset.$
$V((\pi,S)|\nu)\geq\alpha\implies$ $\sum_{s\in S^{1}(\nu)}\pi_{1}(s)=0,\sum_{s\in S^{1}(\nu)}\pi_{n}(s)=1\implies\text{supp}(\pi_{1})\cap\text{supp}(\pi_{n})=\emptyset.$
A.6 Section 7.3.1
Proof of Proposition 3.
Suppose $A$ is ordered. Let $\alpha\in[0,1]$ and let $\beta$ be such that $\mathcal{Q}(\beta)$ is nonempty. Let $\mu^{*}$ be a minimizer for $\Pi(\mu)$ on $\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$, noting (i) $\mathcal{P}(\alpha)\cap\mathcal{Q}(\beta)$ is the intersection of compact sets and therefore compact and (ii) $\Pi$ is continuous by the maximum theorem. Write $(i^{*},z^{*})\equiv(i(\mu^{*}),z(\mu^{*})),(\pi^{*},S^{*})\equiv(\pi^{\mu^{*}},S^{\mu^{*}})$ and choose $d\in\mathcal{D}$.
Consider movements in the positive direction $\mu^{*}+d$. Because $A$ is ordered, we have
$z^{*}d_{i^{*}}+\sum_{j>i^{*}}d_{j}$ $\geq\min(\sum_{j\geq i^{*}}d_{j},\sum_{j>i^{*}}d_{j})\geq 0,$
$\Delta_{i^{*}}z^{*}d_{i^{*}}+\sum_{j>i^{*}}\Delta_{j}d_{j}$ $\geq\min(\sum_{j\geq i^{*}}\Delta_{j}d_{j},\sum_{j>i^{*}}\Delta_{j}d_{j})\geq 0.$
Accordingly, $V((\pi^{*},S^{*})|\mu^{*}+d)\geq V((\pi^{*},S^{*})|\mu^{*})$.
Consider instead movements in the negative direction $\nu^{*}\equiv\mu^{*}-d$. There are three cases to consider. First, if $\nu^{*}$ is ineligible then
$\sum_{i}\Delta_{i}\mu^{*}=\sum_{i}\Delta_{i}\nu_{i}^{*}+\sum_{i}\Delta_{i}d_{i}\geq\sum_{i}\Delta_{i}d_{i}\geq 0$
and thus $\mu^{*}$ is also ineligible. In turn, we have $(i(\mu^{*}),z(\mu^{*}),\Pi(\mu^{*}))=(i(\nu^{*}),z(\nu^{*}),\Pi(\nu^{*}))=(1,1,1)$ and hence $V((\pi^{*},S^{*})|\nu^{*})=V((\pi^{*},S^{*})|\mu^{*})=1$.
Second, if $\nu^{*}$ is eligible and $\mu^{*}$ is not then $\Pi(\mu^{*})=1$ and $\Pi(\nu^{*})<1$. Because $\mu^{*}$ is a minimizer, we conclude $\nu^{*}\notin\mathcal{P}(\alpha)$.
Third, if $\nu^{*}$ and $\mu^{*}$ are both eligible then $\sum_{j\geq i^{*}}\Delta_{j}\nu^{*}\leq\sum_{j\geq i^{*}}\Delta_{j}\mu_{j}^{*}<0,$ where the first inequality follows from the fact $A$ is ordered and the second from the definition of $i^{*}=i(\mu^{*})$. In turn, it follows from the definition of $i(\cdot)$ that $i(\nu^{*})\geq i^{*}$.
If $i(\nu^{*})=i^{*}$ then by the definition of $\Pi(\cdot)$
$\Pi(\mu^{*})-\Pi(\nu^{*})=\sum_{j>i^{*}}d_{j}+\frac{1}{|\Delta_{i^{*}}|}\sum_{j>i^{*}}\Delta_{j}d_{j}.$
Accordingly, if $\max(\sum_{j>i^{*}}d_{j},\sum_{j>i^{*}}\Delta_{j}d_{j})>0$ then $\Pi(\nu^{*})<\Pi(\mu^{*})$ and hence $\nu^{*}\notin\mathcal{P}(\alpha)$. Alternatively, if $\sum_{j>i^{*}}d_{j}=\sum_{j>i^{*}}\Delta_{j}d_{j}=0$ then $d_{i^{*}}=0$ because $\Delta_{i^{*}}<0$ and $\sum_{j\geq i^{*}}d_{j},\sum_{j\geq i^{*}}\Delta_{j}d_{j}\geq 0$. In that case, we apparently have $V((\pi^{*},S^{*})|\nu^{*})=V((\pi^{*},S^{*})|\mu^{*})$.
Otherwise, if $j^{*}\equiv i(\nu^{*})>i^{*}$ then
$\Pi(\nu^{*})$ $=\nu_{j^{*}}^{*}z(\nu^{*})+\sum_{j>j^{*}}\nu^{*}<\sum_{j\geq j^{*}}\nu_{j}^{*}\leq\sum_{j\geq j^{*}}\mu_{j}^{*}\leq\mu_{i^{*}}^{*}z^{*}+\sum_{j>i^{*}}\mu_{j}^{*}=\Pi(\mu^{*}),$
where the strict inequality follows from $\nu_{j^{*}}^{*}z(\nu^{*})<v_{j^{*}}^{*}$ via $j^{*}\in\text{supp}(\nu^{*})$ and $z(\nu^{*})<1$; the first weak inequality follows from our hypothesis that $A$ is ordered; and the second weak inequality follows from our hypothesis that $j^{*}>i^{*}$. We again conclude that $\nu^{*}\notin\mathcal{P}(\alpha)$. ∎
Proof of Lemma 3.
Suppose $A$ is not ordered. By definition, there exists a state $11\leq i\leq L+$ 1 and a direction $d\in\mathcal{D}$ such that at least one of the quantities $\sum_{j\geq i}d_{j},$ $\sum_{j\geq i}d_{j}\Delta_{j}$ is strictly negative. If $i=L+1$ then our normalizations $\mathcal{D}$ imply directly that $\sum_{j\in H}d_{j}>0>\sum_{j\in H}d_{j}\Delta_{j}$. Alternatively, suppose $\sum_{j\in H}d_{j},\sum_{j\in H}d_{j}\Delta_{j}\geq 0$ and let $i$ be the largest state in $L$ with the property that either $\sum_{j\geq i}d_{j}<0$ or $\sum_{j\geq i}d_{j}\Delta_{j}<0$. Because $\sum_{j>i}d_{j},\sum_{j>i}d_{j}\Delta_{j}\geq 0$, we have
$\sum_{j\geq i}d_{j}<0\leq\sum_{j>i}d_{j}\implies d_{i}<0\implies d_{i}\Delta_{i}>0\implies\sum_{j\geq i}d_{j}\Delta_{j}>\sum_{j>i}d_{j}\Delta_{j}\geq 0,$
$\sum_{j\geq i}d_{j}\Delta_{j}<0\leq\sum_{j\geq i}d_{j}\Delta_{j}\implies d_{i}\Delta_{i}<0\implies d_{i}>0\implies\sum_{j\geq i}d_{j}>\sum_{j>i}d_{j}\geq 0.$
If $i>1$, there is nothing more to show. Otherwise, if $i=1$ then $\sum_{j\geq i}d_{j}=\sum_{j\geq 1}d_{j}=0$ and hence $\sum_{j\geq i}d_{j}\Delta_{j}<0$ by hypothesis. Because
$\sum_{j\geq 1}d_{j}\Delta_{j}<0\leq\sum_{j>1}d_{j}\Delta_{j}\implies d_{1}\Delta_{1}<0\implies d_{1}>0\implies\sum_{j\geq 2}d_{j}<0,$
this contradicts our choice of $i$ to satisfy $\sum_{j>i}d_{j}\geq 0$ and we conclude $i>1$. Because we have already established the claim for that case, this completes the proof. ∎
Proof of Lemma 4.
Let $i\in L$ and $z\in(0,1)$. Define $\Delta\equiv\sum_{j>i}\Delta_{j}$ and let $\lambda\in[0,1)$ satisfy $(1-\lambda)\Delta+\lambda(n-1)\Delta_{n}>0$. Further, define constants
$\kappa$ $\equiv\frac{1}{|\Delta_{i}|}((1-\lambda)\Delta+\lambda(n-1)\Delta_{n}),$ $x$ $\equiv\frac{z}{(n-1)z+\kappa}$
and consider prior $\mu$ with $\mu_{i}\equiv 1-(n-1)x,\mu_{n}\equiv(1-\lambda)x+\lambda(n-1)x,$ and $\mu_{j}\equiv(1-\lambda)x$ for all $j\neq i,n$. First, $\mu$ has full support because $\lambda<1$. Second, direct substitution yields $\Delta_{i}z\mu_{i}+\sum_{j>i}\Delta_{j}\mu_{j}=\Delta_{i}\kappa x+|\Delta_{i}|\kappa x=0$ and thus $(i(\mu),z(\mu))=(i,z)$. ∎
Proof of Lemma 5.
Suppose prior $\mu$ has full support. If $z(\mu)\in(0,1)$ then by definition $\sum_{i\geq i(\mu)}\Delta_{i}\mu_{i}<0<\sum_{i>i(\mu)}\Delta_{i}\mu_{i}$ and thus $i(\mu+\varepsilon d)=i(\mu-\varepsilon d)=i(\mu)$. In turn,
$\Pi(\mu+\varepsilon d)$ $=\sum_{i>i(\mu)}(\mu_{i}+\varepsilon d_{i})(1+\frac{\Delta_{i}}{|\Delta_{i(\mu)}|}),$ $\Pi(\mu-\varepsilon d)$ $=\sum_{i>i(\mu)}(\mu_{i}-\varepsilon d_{i})(1+\frac{\Delta_{i}}{|\Delta_{i(\mu)}|})$
and thus $\text{sign}(\Pi(\mu+\varepsilon d)-\Pi(\mu))=-\text{sign}(\Pi(\mu-\varepsilon d))$. If $\Pi(\mu+\varepsilon d)=\Pi(\mu-\varepsilon d)=\Pi(\mu)$ then the proof is complete. Alternatively, if $\Pi(\mu+\varepsilon d)>\Pi(\mu)$ then consider cutoff signal $(\pi^{i,z},S^{i,z})$ with $i\equiv i(\mu)$, $z\equiv(\mu_{i}+\varepsilon d_{i})^{-1}(\Pi(\mu)-\sum_{j>i}(\mu_{j}+\varepsilon d_{j}))$ and note
$\sum_{j}\pi_{j}^{i,z}(1)(\mu_{j}+\varepsilon d_{j})$ $=\Pi(\mu),$ $\sum_{j}\Delta_{j}\pi_{j}^{i,z}(1)(\mu_{j}+\varepsilon d_{j})$ $=\Delta_{j}(\Pi(\mu)-\Pi(\mu+\varepsilon d))>0,$
where the strict inequality follows from $i\in L$ and $\Pi(\mu+\varepsilon d)>\Pi(\mu)$. Accordingly, $V((\pi^{i,z},S^{i,z})|\mu+\varepsilon d)=\Pi(\mu)$ and $\mu+\varepsilon d\in\mathcal{P}(\Pi(\mu))$. The case with $\Pi(\mu-\varepsilon d)>\Pi(\mu)$ is substantively identical. ∎
Proof of Proposition 4.
Suppose $A$ is not ordered. Per Lemma 3, there exists a direction $d\in\mathcal{D}$ and a state $i$ with $11<i<\max L+$ 1 such that one of
$\sum_{j\geq i}\Delta_{j}d_{j}$ $>0>\sum_{j\geq i}d_{j},$ (3)
$\sum_{j\geq i}d_{j}$ $>0>\sum_{j\geq i}\Delta_{j}d_{j}$ (4)
obtains. There are two substantive cases. First, suppose $i\in L$ and consider subcase (3). Choose $z\in(0,1)$ to satisfy $zd_{i}+\sum_{j>i}d_{j}<0<z\Delta_{i}d_{i}+\sum_{j>i}\Delta_{j}d_{i}$ and recall Lemma 4 provides a prior $\mu$ with full support and $(i(\mu),z(\mu))=(i,z)$. In turn, Proposition 2 implies $(\pi,S)$ satisfies $V((\pi,S)|\mu)\geq V((\pi^{\prime},S^{\prime})|\mu)$ for all signals $(\pi^{\prime},S^{\prime})$ if and only if $(\pi,S)$ is equivalent to $(\pi^{\mu},S^{\mu})$. Consider the experimenter’s problem with data $(\alpha,\beta)\equiv(\Pi(\mu),A\mu)$ and let $(\pi,S)$ be any such signal. Per Lemma 5, either $\mu+\varepsilon d\in\mathcal{P}(\alpha)$ or $\mu-\varepsilon d\in\mathcal{P}(\alpha)$. If $\mu+\varepsilon d\in\mathcal{P}(\alpha)$ then $S^{0}(\mu)\subset S^{0}(\mu+\varepsilon d)$ by $\sum_{i}\Delta_{i}(\mu+\varepsilon d)_{i}^{s}\approx\sum_{i}\Delta_{i}\mu_{i}^{s}<0$ for all messages $s\in S^{0}(\mu)$ and thus
$V((\pi,S)|\mu+\varepsilon d)=\sum_{j}\sum_{s\in S^{1}(\mu+\varepsilon d)}\pi_{j}(s)(\mu_{j}+\varepsilon d_{j})\leq\sum_{j}\sum_{s\in S^{1}(\mu)}\pi_{j}(s)(\mu_{j}+\varepsilon d_{j}).$
Because our choice of $z$ implies
$\sum_{j}\sum_{s\in S^{1}(\mu)}\pi_{j}(s)(\mu_{j}+\varepsilon d_{j})$ $=\alpha+\varepsilon(zd_{i}+\sum_{j>i}d_{j})<\alpha,$
we have $V((\pi,S)|\mu+\varepsilon d)<\alpha$. If instead $\mu-\varepsilon d\in\mathcal{P}(\alpha)$, then we have $S^{0}(\mu)\subset S^{0}(\mu-\varepsilon d)$ by the same argument as above and thus
$V((\pi,S)|\mu-\varepsilon d)=\sum_{j}\sum_{s\in S^{1}(\mu-\varepsilon d)}\pi_{j}(s)(\mu_{j}-\varepsilon d_{j})\leq\sum_{j}\sum_{s\in S^{1}(\mu)}\pi_{j}(s)(\mu_{j}-\varepsilon d_{j}).$
Because our choice of $z$ implies
$\sum_{j}\sum_{s\in S^{1}(\mu)}\Delta_{j}\pi_{j}(s)(\mu_{j}-\varepsilon d_{j})$ $=-\varepsilon(z\Delta_{i}d_{i}+\sum_{j>i}\Delta_{j}d_{j})<0,$
there exists a message $s^{*}\in S^{0}(\mu-\varepsilon d)\cap S^{1}(\mu)$ with $\sum_{j}\pi_{j}(s^{*})(\mu_{j}-\varepsilon d_{j})>0$. Together, we obtain
$V((\pi,S)|\mu-\varepsilon d)\leq\sum_{j}\sum_{s\in S^{1}(\mu)\setminus\{s^{*}\}}\pi_{j}(s)(\mu_{j}-\varepsilon d_{j})<\sum_{j}\sum_{s\in S^{1}(\mu)}\pi_{j}(s)(\mu_{j}-\varepsilon d_{j})\approx\alpha.$
This establishes the argument for case $i\in L$ in subcase (3). Subcase (4) is symmetric.
Second, suppose instead $i=\max L+1$ and consider subcase (3). Choose $z\in(0,1)$ to satisfy $zd_{i-1}+\sum_{j\geq i}d_{j}<0<zd_{i-1}\Delta_{i-1}+\sum_{j\geq i}d_{j}\Delta_{j}$ and recall again that Lemma 4 provides a cutoff signal $(i-1,z)$ and prior with full support $\mu$ such that $(i(\mu),z(\mu))=(i-1,z)$. At this point, The argument is the same as in the $i\in L$ case, with subcase (4) again symmetric to its counterpart. ∎
A.7 Section 7.3.2
Proof of Proposition 5.
We write $d^{ij}$ for the vector in $\mathbb{R}^{n}$ satisfying $d_{i}^{ij}\equiv 1$ and $d_{j}^{ij}\equiv-1$ with 00 in all other entries. Suppose $A$ is simple. Let $\{\Sigma_{1},...,\Sigma_{k}\}$ be the associated partition of $\Omega$, let $\tau\equiv(\tau_{1},...,\tau_{k})$ be an arbitrary selection from the correspondence $t\rightrightarrows\Sigma_{t}$, define experiment $B\equiv(B_{1},...,B_{k})$ by $B_{t}\equiv A_{\tau_{t}}$, and let $\mu$ be any prior with full support. First, we claim $A_{i}=B_{t}$ for all $i\in\Sigma_{t}$. To see why, let $i,j\in\Sigma_{t}$ be distinct and define $\nu\equiv\mu+\varepsilon d^{ij}$. Because $A$ is simple and $\nu(\Sigma_{t})=\mu(\Sigma_{t})$ for all $t$, we have $\nu\in\mathcal{Q}(A\mu)$ and hence $(\nu-\mu)\in\mathcal{D}$. Accordingly, $Ad^{ij}=A_{i}-A_{j}=0$, as claimed. Second, we claim $B$ has full rank. To see why, suppose to the contrary that $Bx=0$ for some $(x_{1},...,x_{k})\neq(0,...,0)$ and let $(y_{1},...,y_{n})$ be any vector satisfying $\sum_{i\in\Sigma_{j}}y_{i}=x_{j}$ for all indices $j$. Because $Ay=Bx=0$, we have $\nu\equiv(\mu+\varepsilon y)\in\mathcal{Q}(A\mu)$. Further, because there exists index $t$ with $\sum_{i\in\Sigma_{t}}y_{i}=x_{t}\neq 0$, we have $\nu(\Sigma_{t})\neq\mu(\Sigma_{t})$. This contradicts our hypothesis that $A$ is simple, and we conclude that $B$ indeed has full rank.
Conversely, suppose there exists a partition $\{\Sigma_{1},...,\Sigma_{k}\}$ of $\Omega$ and an experiment $B\equiv(B_{1},...,B_{k})$ such that (i) $A_{i}=B_{t}$ for all $i\in\Sigma_{t}$ and (ii) $B$ has full rank. Let $\mu,\nu\in\Delta(\Omega)$, define $(x_{1},...,x_{k})\equiv(\mu(\Sigma_{1})-\nu(\Sigma_{1}),...,\mu(\Sigma_{k})-\nu(\Sigma_{k}))$, and note that $\nu\in\mathcal{Q}(A\mu)\iff A(\mu-\nu)=Bx=0\iff x=0$. Accordingly, $A$ is simple. ∎
Proof of Theorem 5.
We again write $d^{ij}$ for the vector in $\mathbb{R}^{n}$ satisfying $d_{i}^{ij}\equiv 1$ and $d_{j}^{ij}\equiv-1$ with 00 in all other entries. Our approach is to show that $A$ is ordered if and only if it satisfies one of the two criteria in the statement of the result. Theorem 4 then implies the result.
Suppose $A$ separates every low state and choose $d\in\mathcal{D}$. Because $\sum_{i\in L}d_{i}=\sum_{i\in L}0=0$ by hypothesis and $\sum_{i}d_{i}=0$ by definition, we have $\sum_{i\in H}d_{i}=0$. By normalization, then, we also have $\sum_{i\in H}\Delta_{i}d_{i}\geq 0$. It follows quickly from $d_{i}=0$ for all $i\in L$ that $\sum_{j\geq i}d_{i},\sum_{j\geq i}\Delta_{i}d_{i}\geq 0$ for all $j=1,...,\max L+1$, and we conclude that $A$ is ordered.
Alternatively, suppose that $A$ pools exactly one pair of states $(i,j)\in H\times L$ and separates all other states. It is immediate that (i) $\mathcal{D}=\{\kappa d^{ij}|\kappa\geq 0\}$ and (ii) $\kappa d^{ij}$ satisfies both the frequency and the incentive compatibility criteria for all $\kappa\geq 0$. Accordingly, $A$ is ordered.
Conversely, suppose $A$ satisfies neither property. There are four subcases. First, if there exist states $j<i\leq\max L$ with $i,j\in\Sigma_{t}$, then $d^{ij}\in\mathcal{D}$. Because $\sum_{i^{\prime}\geq i}\Delta_{i^{\prime}}d_{i^{\prime}}<0$, $A$ is not ordered. Second, if there exist distinct states $i\leq\max L<j<k$ then $\mathcal{D}$ contains the direction $d\equiv d^{jk}+\varepsilon d^{ji}$. Because $\sum_{\tau\geq j}\Delta_{\tau}d_{\tau}=(1+\varepsilon)\Delta_{j}-\Delta_{k}<0$, $A$ is not ordered. Third, if there exist states $i<j\leq\max L<k<l$ and indices $t,t^{\prime}$ such that $\Sigma_{t}=\{i,k\},\Sigma_{t^{\prime}}=\{j,l\}$, then $\mathcal{D}$ contains the direction $d\equiv d^{jl}+(1+\varepsilon)d^{ki}$. Because $\sum_{\tau\in H}\Delta_{\tau}d_{\tau}=(1+\varepsilon)\Delta_{k}-\Delta_{l}<0$, $A$ is not ordered. Fourth, and finally, if there exist states $i,j,k,l$ with $i\leq\max L<j,$ $\max L<k<l$ and indices $t,t^{\prime}$ with $\Sigma_{t}=\{i,j\}$ and $k,l\in\Sigma_{t^{\prime}}$ then $\mathcal{D}$ contains the direction $d\equiv d^{sk}+\varepsilon d^{ji}$. Because $\sum_{\tau\in H}\Delta_{\tau}d_{\tau}=\varepsilon\Delta_{j}+(\Delta_{k}-\Delta_{l})<0$, $A$ is not ordered. ∎
A.8 Section 7.3.3
Proof of Lemma 6.
Write $i^{0}$ for the state $i$ with $\Delta_{i^{0}}=0$ and recall that Lemma 4 provides a fully supported prior $\mu$ with $(i(\mu),z(\mu))=(i,z)$, $\Delta_{i}\mu_{i}z+\sum_{j>i}\Delta_{j}\mu_{j}=0$. Because $\mu$ has full support and $z>0$, we also have $\Pi(\mu)>0$. If $\Pi(\mu)\leq\alpha$, define
$\nu\equiv\frac{1-\alpha}{1-\Pi(\mu)}\mu+\frac{\alpha-\Pi(\mu)}{1-\Pi(\mu)}\delta(\omega_{i^{0}}).$
Alternatively, if $\Pi(\mu)>\alpha$ define
$\nu\equiv\frac{\alpha}{\Pi(\mu)}\delta(\omega_{1})+\frac{\Pi(\mu)-\alpha}{\Pi(\mu)}\mu.$
In both cases, we have $\nu_{i}z+\sum_{j>i}\nu=\alpha$ and $\Delta_{i}\nu_{i}z+\sum_{j>i}\Delta_{j}\nu_{j}=0.$ Because $\nu$ inherits full support from $\mu$, it meets all the criteria of the claim. ∎
Proof of Theorem 6.
Fix action distribution $\alpha$. First, if $A$ is ordered then Theorem 4 implies $A$ is reliable. Because reliability implies sequential reliability, $A$ is sequentially reliable. Conversely, if $A$ is not ordered then the non-reliability of $A$ follows from Lemma 3, Lemma 5, and Lemma 6 from the same argument given in the proof of Proposition 4, with Lemma 6 as a surrogate for Lemma 4. ∎